# ART: Adaptive Relation Tuning for Generalized Relation Prediction

### v1

@@ 2 Related work
Uncertainty estimation is essential in setups that require informative sample selection. At a high level, uncertainty can be classified into aleatoric (data) and epistemic (model) uncertainty. Aleatoric uncertainty reflects intrinsic variability in the data, while epistemic uncertainty arises from the model’s limited understanding of the underlying task. Some of the commonly used uncertainty estimators are Monte-Carlo dropout [8], deep ensembles [7], and the entropy of the softmax predictions [27]. These uncertainty measures are used predominately to compute model uncertainty. In our work, we use entropy as an uncertainty measure due to its cost-effective use in a VLM setup compared to dropout and ensembles. We also use a cosine similarity-based approach [18] as an additional uncertainty measure.
@@ 3 Adaptive relation tuning framework
VRD models must generalize well to be effective in real-world scenarios. Yet, models trained solely on VRD data-sets either overfit the underlying distribution or learn spurious correlations. Though VLMs excel in generalization due to their large training corpora, they often struggle with fine-grained relations, favoring frequent patterns. To address this, ART follows a two-step process: (1) Relation-tuning data creation, where the training data is structured with high-level relation categories (spatial, semantic, possessive) and a counter-negative mining strategy to refine negative samples; and (2) Adaptive relation tuning of VLMs, which optimizes the model to focus on rare, diverse relations, improving generalization to unseen scenarios.
@@ 3.1 Relation-tuning data creation
A key aspect of our approach is relation-tuning data construction, which involve carefully crafting questions to capture fine-grained object relations. A simple transformation of the ground-truth relation triplets $\langle s,p,o\rangle$ into a question format such as “Is there a relation between $\langle s\rangle$ and $\langle o\rangle$ in the image?”—with the answers “Yes, $\langle s,p,o\rangle$” or “No” for valid or invalid relations, respectively, might seem sufficient. However, this naive approach overlooks important nuances in relation classification, as shown below (Tab. 2 in Sec. 4.2). To address this, we introduce a more structured approach to relation-tuning data creation, where we design a question format that incorporates two key components: (1) high-level relation categories and (2) counter-negative mining. The final relation tuning format template, shown in Fig. 2, is derived from this structured formulation, which is empirically supported by findings in Tab. 2.
@@
[…]
@@ 3.2 Adaptive relation tuning (ART)
Step 1: Sampling budget allocation. To prioritize predicates requiring greater supervision, we compute a per-predicate sampling probability as $P_{p}=\frac{1-R_{p}}{\sum_{j}(1-R_{j})}$, where $R_{p}$ is the recall of predicate $p$ on the validation set. Each predicate is allocated budget $B^{\prime}_{p}=\min(B\cdot P_{p},N_{p})$, ensuring that predicates with lower recall receive a higher sampling budget while preventing over-sampling beyond available instances (Lines 7–10).
@@
$\mathcal{R}\leftarrow$ Recall per predicate on validation set $\mathcal{S}_{\text{val}}$
@@
9:  $P_{p}\leftarrow\frac{1-R_{p}}{\sum_{j}(1-R_{j})}$triangleright Higher weight for lower recall
10:  $B^{\prime}_{p}\leftarrow\min(B\cdot P_{p},N_{p})$triangleright $N_{p}$: available data samples for $p$
@@
32:     $z\leftarrow z-0.1$triangleright adjust threshold if insufficient samples
@@ 4.1 Experimental setup
Evaluation protocol and metrics.  As our aim is to improve predicate prediction in a generalizable manner, we evaluate model performance on the predicate classification task using ground-truth/predicted bounding boxes and objects, with a focus on generalization across data distributions, object, and predicate categories. Following previous works [44, 34], we report Recall@k (R@k) and mean Recall (mR@k). Mean recall (mR@k) is particularly important as it reflects a model’s ability to perform across the full predicate distribution. To further assess generalization, we propose generalized Recall (gR@k) and mean generalized Recall (mgR@k), which treat false positives with high semantic similarity (measured by $\mathbf{S}$ from Sec. 3.2) as true positives. See Appendix B for details.
[…]
@@ 4.2 Experimental results
LISA [17] 7.877.877.87 18.9218.9218.92
DeiSAM + GT (w/o bbox) 4.774.774.77 11.8411.8411.84
DeiSAM + ART (w/o bbox) 25.0725.0725.07 26.6226.6226.62
DeiSAM + GT (with bbox) 35.0435.0435.04 43.8443.8443.84
DeiSAM + ART (with bbox) 99.9899.9899.98 97.9697.9697.96
@@ 5 Conclusion
In this work, we presented ART, a novel framework for the relation classification task of VRD that leverages vision-language models through instruction tuning by converting benchmark VRD datasets into an instruction-tuning format. We ensured that our model focuses on the most relevant instances by incorporating adaptive sampling and fine-tuning techniques, thus enhancing its robustness and generalization capabilities. ART consistently outperformed baseline models, improving the handling of both diverse and unseen object classes and relations. This underscores the importance of innovative VRD approaches, especially for real-world applications with diverse data distributions. Future research can enhance ART by refining instruction tuning strategies and extending it to other vision-language tasks, potentially unlocking new applications in visual understanding.
@@ B Additional experimental details
In addition to the hyperparameters outlined in Sec. 4 of the main paper, we set the initial $z$-score threshold to 1.96, which corresponds to 95% of the data. This threshold was chosen because a $z$-score of 1.96 is more sensitive to potential outliers, making it useful for capturing more subtle deviations from the norm. For training, we use an initial learning rate of $1e-3$ and a linear warmup for 3000 steps. We optimize with Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and apply a weight decay of 0.05. Additionally, we chose 12% of training data for instruction tuning as mR@k saturates near this point as analyzed in Fig. 5. We use the LAVIS library [19] for implementation, training, and evaluation. The models are trained using four Nvidia A100 (40Gb) GPUs within two days.
@@ C Balanced sampling
2:   $\mathcal{P}=\{p_{i}\}_{i=1}^{N}$triangleright Set of $N$ predicate categories
3:   $\mathcal{S}_{\text{train}}\leftarrow\emptyset$triangleright Initial training set
4:   $\mathcal{S}_{\text{pool}}$triangleright Remaining dataset (excluding $\mathcal{S}_{\text{train}}$)
@@
9:   $B^{\prime}_{p}\leftarrow 0,\quad\forall p\in\mathcal{P}$triangleright Initialize per-predicate budget
@@
13:   $B^{\prime}_{p_{i}}\leftarrow B^{\prime}_{p_{i}}+1$triangleright Allocate one sample to predicate $p_{i}$
14:   $N_{p_{i}}\leftarrow N_{p_{i}}-1$triangleright Decrement available samples
15:   $B\leftarrow B-1$triangleright Reduce remaining budget
@@
17:  $i\leftarrow(i+1)\mod N$triangleright Move to the next predicate
@@ F.3 On R@k and gR@k performance trade-offs
[…]

### v2

@@ 2 Related work
Uncertainty estimation is essential in setups that require informative sample selection. At a high level, uncertainty can be classified into aleatoric (data) and epistemic (model) uncertainty. Aleatoric uncertainty reflects intrinsic variability in the data, while epistemic uncertainty arises from the model’s limited understanding of the underlying task. Some of the commonly used uncertainty estimators are Monte-Carlo dropout [8], deep ensembles [7], and the entropy of the softmax predictions [27]. These uncertainty measures are used predominantly to compute model uncertainty. In our work, we use entropy as an uncertainty measure due to its cost-effective use in a VLM setup compared to dropout and ensembles. We also use a cosine similarity-based approach [18] as an additional uncertainty measure.
@@ 3 Adaptive relation tuning framework
VRD models must generalize well to be effective in real-world scenarios. Yet, models trained solely on VRD datasets either overfit the underlying distribution or learn spurious correlations. Though VLMs excel in generalization due to their large training corpora, they often struggle with fine-grained relations, favoring frequent patterns. To address this, ART follows a two-step process: (1) Relation-tuning data creation, where the training data is structured with high-level relation categories (spatial, semantic, possessive) and a counter-negative mining strategy to refine negative samples; and (2) Adaptive relation tuning of VLMs, which optimizes the model to focus on rare, diverse relations, improving generalization to unseen scenarios.
@@ 3.1 Relation-tuning data creation
A key aspect of our approach is relation-tuning data construction, which involves carefully crafting questions to capture fine-grained object relations. A simple transformation of the ground-truth relation triplets $\langle s,p,o\rangle$ into a question format such as “Is there a relation between $\langle s\rangle$ and $\langle o\rangle$ in the image?”—with the answers “Yes, $\langle s,p,o\rangle$” or “No” for valid or invalid relations, respectively, might seem sufficient. However, this naive approach overlooks important nuances in relation classification, as shown below (Tab. 2 in Sec. 4.2). To address this, we introduce a more structured approach to relation-tuning data creation, where we design a question format that incorporates two key components: (1) high-level relation categories and (2) counter-negative mining. The final relation tuning format template, shown in Fig. 2, is derived from this structured formulation, which is empirically supported by findings in Tab. 2.
@@
$\text{positive: }r$ $\in\{\text{spatial},\text{possessive},\text{semantic}\},$ (1)
$\text{negative: }nr$ $=\begin{cases}\{\text{possessive}\},&\text{if }r=\text{spatial},\\ \{\text{spatial},\text{semantic}\},&\text{if }r=\text{possessive},\\ \{\text{possessive}\},&\text{if }r=\text{semantic},\end{cases}$
@@ 3.2 Adaptive relation tuning (ART)
Step 1: Sampling budget allocation. To prioritize predicates requiring greater supervision, we compute a per-predicate sampling probability as $P_{p}=\frac{1-R_{p}}{\sum_{j}(1-R_{j})}$, where $R_{p}$ is the recall of predicate $p$ on the validation set. Each predicate is allocated a budget $B^{\prime}_{p}=\min(B\cdot P_{p},N_{p})$, ensuring that predicates with lower recall receive a higher sampling budget while preventing over-sampling beyond available instances (Lines 7–10).
@@
[…]
@@
9:  $P_{p}\leftarrow\frac{1-R_{p}}{\sum_{j}(1-R_{j})}$ $\triangleright$ Higher weight for lower recall
10:  $B^{\prime}_{p}\leftarrow\min(B\cdot P_{p},N_{p})$ $\triangleright$ $N_{p}$: available data samples for $p$
@@
32:     $z\leftarrow z-0.1$ $\triangleright$ adjust threshold if insufficient samples
@@ 4.1 Experimental setup
Evaluation protocol and metrics.  As our aim is to improve predicate prediction in a generalizable manner, we evaluate model performance on the predicate classification task using ground-truth/predicted bounding boxes and objects, with a focus on generalization across data distributions, objects, and predicate categories. Following previous works [44, 34], we report Recall@k (R@k) and mean Recall (mR@k). Mean recall (mR@k) is particularly important as it reflects a model’s ability to perform across the full predicate distribution. To further assess generalization, we propose generalized Recall (gR@k) and mean generalized Recall (mgR@k), which treat false positives with high semantic similarity (measured by $\mathbf{S}$ from Sec. 3.2) as true positives. See Appendix B for details.
Implementation details.  We leverage BLIP-2 [20] models fine-tuned for the captioning task with strong generalization capabilities, specifically two InstructBLIP [5] variants. Both share the ViT/14 [6] image encoder but differ in language models: One integrates a Vicuna-7B [49] LLM, instruction-tuned from LLaMA [37], while the other uses FlanT5 [4], based on Transformer T5 [30]. We fine-tune both variants with our adaptive instruction tuning. The models reached the best results with about $12\%$ (see Appendix B) of the training data samples, with a sampling budget of $2\%$ per adaptive learning loop.
@@ 4.2 Experimental results
LISA [17] $7.87$ $18.92$
DeiSAM + GT (w/o bbox) $4.77$ $11.84$
DeiSAM + ART (w/o bbox) $25.07$ $26.62$
DeiSAM + GT (with bbox) $35.04$ $43.84$
DeiSAM + ART (with bbox) $99.98$ $97.96$
@@ 5 Conclusion
In this work, we presented ART, a novel framework for the relation classification task of VRD that leverages vision-language models through instruction tuning by converting benchmark VRD datasets into an instruction-tuning format. Our model focuses on the most relevant instances by incorporating adaptive sampling and fine-tuning techniques, thus enhancing its robustness and generalization capabilities. ART consistently outperforms baseline models, improving the handling of both diverse and unseen object classes and relations. This underscores the importance of innovative VRD approaches, especially for real-world applications with diverse data distributions. Future research can enhance ART by refining instruction tuning strategies and extending it to other vision-language tasks, potentially unlocking new applications in visual understanding.
@@ B Additional experimental details
In addition to the hyperparameters outlined in Sec. 4 of the main paper, we set the initial $z$-score threshold to 1.96, which corresponds to 95% of the data. This threshold was chosen because a $z$-score of 1.96 is more sensitive to potential outliers, making it useful for capturing more subtle deviations from the norm. For training, we use an initial learning rate of 1e31e-3 and a linear warmup for 3000 steps. We optimize with Adam ($\beta_{1}=0.9$, $\beta_{2}=0.999$) and apply a weight decay of 0.05. Additionally, we chose 12% of training data for instruction tuning as mR@k saturates near this point as analyzed in Fig. 5. We use the LAVIS library [19] for implementation, training, and evaluation. The models are trained using four Nvidia A100 (40Gb) GPUs within two days.
@@ C Balanced sampling
2:   $\mathcal{P}=\{p_{i}\}_{i=1}^{N}$ $\triangleright$ Set of $N$ predicate categories
3:   $\mathcal{S}_{\text{train}}\leftarrow\emptyset$ $\triangleright$ Initial training set
4:   $\mathcal{S}_{\text{pool}}$ $\triangleright$ Remaining dataset (excluding $\mathcal{S}_{\text{train}}$)
@@
9:   $B^{\prime}_{p}\leftarrow 0,\quad\forall p\in\mathcal{P}$ $\triangleright$ Initialize per-predicate budget
@@
13:   $B^{\prime}_{p_{i}}\leftarrow B^{\prime}_{p_{i}}+1$ $\triangleright$ Allocate one sample to predicate $p_{i}$
14:   $N_{p_{i}}\leftarrow N_{p_{i}}-1$ $\triangleright$ Decrement available samples
15:   $B\leftarrow B-1$ $\triangleright$ Reduce remaining budget
@@
17:  $i\leftarrow(i+1)\mod N$ $\triangleright$ Move to the next predicate
@@ F.3 On R@k and gR@k performance trade-offs
Table 7: Adaptive vs. fixed thresholding. $t_{\text{FP}}$: low-similarity FP threshold, $t_{\text{FN}}$: low-entropy FN threshold, $h_{\text{FN}}$: high-entropy FN threshold, $h_{\text{TP}}$: high-entropy TP threshold. From the mid point threshold, we increase (higher-$h$) or decrease (lower-$t$) the respective thresholds to analyse the effect of fixed thresholds.